# Two-stage Framework for Robust Speech Emotion Recognition Using Target Speaker Extraction in Human Speech Noise Conditions

### v1

@@ 1 Introduction
Speech emotion recognition (SER), as a branch of affective computing, has garnered growing attention over the past two decades because of its contribution to human-computer interactions [Schuller18_Review]. Generally, the mechanism of SER involves extracting and classifying effective emotional features from audio signals so that various emotions of a speaker can be captured, thanks to which SER has been applied in healthcare [Uddin20_healthcare, Hossain15_healthcare], driver safety [Grimm07_driver, Tawari10_driver], call center [Petrushin99_call, Morrison07_call], and online education [Li07_education, Tickle13_education]. At present, research on SER systems on scenarios devoid of background noises, often referred to as clean scenarios, has shown good performance [Liu23_Discriminative, Shen23_Temporalshift, Chen23_Exploring]. However, in real-world environments, the performance of SER significantly degrades, mainly due to the presence of various noises from different sources. These unknown noises severely affect the performance of SER systems, which poses major challenges for the widespread application of SER systems.
@@ 3.1 Datasets
IEMOCAP: The Interactive Emotional Dyadic Motion Capture (IEMOCAP) corpus [Busso08_IEMOCAP], consisting of approximately 12 hours of recordings, includes five dyadic sessions, each with one English male speaker and one English female speaker. For our experiments, we used IEMOCAP as the clean emotional dataset, where we considered only four emotional categories of happy, angry, sad, and neutral. Note that “excited” was merged with “happy” to ensure category balance [Chen23_Exploring, Zen22_CoAttention, guo2021representation].
@@ 4 Conclusion
In this work, we presented a two-stage framework to mitigate the impact of human speech noise on SER. Based on the framework, we designed two training methods, TSE-SER-base and TSE-SER-ft. The effectiveness and robustness of both methods have been verified. Moreover, we investigated the impact of human speech noise on SER, especially on same- and different-gender mixtures. In the future, we plan to explore how our proposed framework performs on human speech noise with different attributes, such as emotion classes and languages. Another possible direction involves multi-interference environments, such as the noisy and reverberant environments.

### v2

@@ 1 Introduction
Speech emotion recognition (SER), as a branch of affective computing, has garnered growing attention over the past two decades because of its contribution to human-computer interactions [Schuller18_Review, shi2020dimensional]. Generally, the mechanism of SER involves extracting and classifying effective emotional features from audio signals so that various emotions of a speaker can be captured, thanks to which SER has been applied in healthcare [Uddin20_healthcare, Hossain15_healthcare], driver safety [Grimm07_driver, Tawari10_driver], call center [Petrushin99_call, Morrison07_call], and online education [Li07_education, Tickle13_education]. At present, research on SER systems on scenarios devoid of background noises, often referred to as clean scenarios, has shown good performance [Liu23_Discriminative, Shen23_Temporalshift, Chen23_Exploring]. However, in real-world environments, the performance of SER significantly degrades, mainly due to the presence of various noises from different sources. These unknown noises severely affect the performance of SER systems, which poses major challenges for the widespread application of SER systems.
@@ 3.1 Datasets
IEMOCAP: The Interactive Emotional Dyadic Motion Capture (IEMOCAP) corpus [Busso08_IEMOCAP], consisting of approximately 12 hours of recordings, includes five dyadic sessions, each with one English male speaker and one English female speaker. For our experiments, we used IEMOCAP as the clean emotional dataset, where we considered only four emotional categories of happy, angry, sad, and neutral. Note that “excited” was merged with “happy” to ensure category balance [Chen23_Exploring, Zen22_CoAttention, guo2021representation, shi2023emotion, shi2024Multimodal].
@@ 4 Conclusion
In this work, we presented a two-stage framework to mitigate the impact of human speech noise on SER. Based on the framework, we designed two training methods, TSE-SER-base and TSE-SER-ft. The effectiveness and robustness of both methods have been verified. Moreover, we investigated the impact of human speech noise on SER, especially on same- and different-gender mixtures. In the future, we plan to explore how our proposed framework performs on human speech noise with different attributes, such as emotion classes and languages. Another possible direction involves multi-interference environments, such as noisy and reverberant environments.